# EchoJEPA: A Latent Predictive Foundation Model for Echocardiography

### v2

@@ Abstract
Foundation models for echocardiography often struggle to disentangle anatomical signal from the stochastic speckle and acquisition artifacts inherent to ultrasound. We present EchoJEPA, a foundation model trained on 18 million echocardiograms across 300K patients, representing the largest pretraining corpus for this modality to date. By leveraging a latent predictive objective, EchoJEPA learns robust anatomical representations that ignore speckle noise. We validate this using a novel multi-view probing framework with frozen backbones, where EchoJEPA outperforms state-of-the-art baselines by approximately 20% in left ventricular ejection fraction (LVEF) estimation and 17% in right ventricular systolic pressure (RVSP) estimation. The model also exhibits remarkable sample efficiency, reaching 79% view classification accuracy with only 1% of labeled data versus 42% for the best baseline trained on 100%. Crucially, EchoJEPA demonstrates superior generalization, degrading by only 2% under physics-informed acoustic perturbations compared to 17% for competitors. Most remarkably, its zero-shot performance on pediatric patients surpasses fully fine-tuned baselines, establishing latent prediction as a superior paradigm for robust, generalizable medical AI.
@@ 3.5 Robustness Evaluation Protocol
Depth attenuation.
Ultrasound signal intensity decreases with tissue depth due to absorption and scattering, particularly in patients with obesity or poor acoustic windows. We simulate this by applying a linear intensity ramp:
@@
where $H$ is image height, $y$ is vertical position, and $\alpha\in\{0.3,0.5,0.7\}$ controls severity.
Acoustic shadow.
Shadows occur when ultrasound is blocked by highly reflective structures such as ribs or calcifications. We simulate this with a Gaussian-weighted intensity reduction:
@@ 4.1 Experimental Setup
Table 3: LVEF estimation (MAE %, lower is better). EchoJEPA-G achieves state-of-the-art across sites using frozen probes, surpassing end-to-end fine-tuned baselines.
@@ 4.3 Sample Efficiency
Table 4: Sample efficiency on view classification (Accuracy %). The EchoJEPA models at 1% labels match or exceed baselines trained on 100% of the data. Values represent mean $\pm$ standard deviation over 3 runs.
@@ 4.5 Multi-View Physiological Estimation
Table 6: Multi-view RVSP estimation (MAE mmHg). Results validate the benefit of early fusion with stream embeddings.
@@ 4.6 Generalization: Adult to Pediatric Transfer
Table 7: Adult to pediatric (LVEF MAE, lower is better). EchoJEPA-G zero-shot outperforms baselines after fine-tuning.
@@ 5 Discussion
EchoJEPA matching state-of-the-art performance with only 10% of labeled data addresses a primary bottleneck in medical AI. Achieving these results with frozen backbones lowers the computational barrier, enabling clinical researchers to train lightweight probes without end-to-end fine-tuning.
@@
EchoJEPA opens directions including fine-grained tasks such as valve segmentation, prospective validation on patients with poor acoustic windows, integration with interpretable reasoning frameworks for clinical decision support (Fallahpour et al., 2025), and extension to other noisy modalities such as fetal and lung ultrasound.
@@ 6 Conclusion
We introduce EchoJEPA, a foundation model for echocardiography demonstrating that latent prediction outperforms pixel reconstruction and alternative pretraining objectives. Trained on 18 million videos across 300K patients, EchoJEPA achieves state-of-the-art performance on LVEF estimation, RVSP prediction, and view classification, reducing LVEF error by 27% over compute-matched reconstruction baselines. EchoJEPA reaches 78.6% view accuracy with 1% of labels versus 42.1% for the best baseline at 100%, degrades by only 2.3% under acoustic perturbations compared to 16.8% for the next-best foundation model, and transfers zero-shot to pediatric patients better than fine-tuned baselines. We release EchoJEPA-L trained on public MIMIC-IV-Echo data alongside our multi-view probing framework.

### v3

@@ Abstract
Foundation models for echocardiography often struggle to disentangle anatomical signal from the stochastic speckle and acquisition artifacts inherent to ultrasound. We present EchoJEPA, a foundation model trained on 18 million echocardiograms across 300K patients, representing the largest pretraining corpus for this modality to date. By leveraging a latent predictive objective, EchoJEPA learns robust anatomical representations that ignore speckle noise. We validate this using a novel multi-view probing framework with frozen backbones, where EchoJEPA outperforms leading baselines by approximately 20% in left ventricular ejection fraction (LVEF) estimation and 17% in right ventricular systolic pressure (RVSP) estimation. The model also exhibits remarkable sample efficiency, reaching 79% view classification accuracy with only 1% of labeled data versus 42% for the best baseline trained on 100%. Crucially, EchoJEPA demonstrates superior generalization, degrading by only 2% under physics-informed acoustic perturbations compared to 17% for competitors. Most remarkably, its zero-shot performance on pediatric patients surpasses fully fine-tuned baselines, establishing latent prediction as a superior paradigm for robust, generalizable medical AI.
@@ 3.5 Robustness Evaluation Protocol
Depth attenuation. Ultrasound signal intensity decreases with tissue depth due to absorption and scattering, particularly in patients with obesity or poor acoustic windows. We simulate this by applying a linear intensity ramp:
@@
where $I(x,y)$ is the original pixel intensity, $I^{\prime}(x,y)$ is the perturbed intensity, $x$ and $y$ are horizontal and vertical pixel coordinates, $H$ is image height, and $\alpha\in\{0.3,0.5,0.7\}$ controls attenuation severity.
Acoustic shadow. Shadows occur when ultrasound is blocked by highly reflective structures such as ribs or calcifications. We simulate this with a Gaussian-weighted intensity reduction along the horizontal axis:
@@ 4.1 Experimental Setup
Table 3: LVEF estimation (MAE, lower is better). All models evaluated with frozen backbones and identical attentive probes. Probes trained on Toronto; Chicago results demonstrate cross-site generalization.
@@ 4.3 Sample Efficiency
Table 4: Sample efficiency on view classification (Accuracy %). EchoJEPA at 1% labels outperforms all baselines trained on 100%. Values represent mean $\pm$ standard deviation over 3 runs.
@@ 4.5 Multi-View Physiological Estimation
Table 6: Multi-view RVSP estimation (MAE mmHg, lower is better). Predictions evaluated with multi-view probing framework.
@@ 4.6 Generalization: Adult to Pediatric Transfer
Table 7: Adult to pediatric transfer (LVEF MAE, lower is better). Zero-shot: probes trained on adult data (EchoNet-Dynamic), evaluated on pediatric data (EchoNet-Pediatric). Fine-tuned: probes trained on EchoNet-Pediatric. EchoJEPA-G zero-shot outperforms all baselines even after probe fine-tuning on pediatric data.
@@ 5 Discussion
EchoJEPA outperforming baselines with only 1% of labeled data addresses a primary bottleneck in medical AI. Achieving strong performance with frozen backbones lowers the computational barrier, enabling clinical researchers to train lightweight probes without end-to-end fine-tuning.
@@
EchoJEPA opens directions including fine-grained tasks such as valve segmentation, prospective validation on patients with poor acoustic windows, integration with interpretable reasoning frameworks for clinical decision support, cardiac dynamics modeling for treatment effect prediction, and extension to other noisy modalities such as fetal and lung ultrasound.
@@ 6 Conclusion
We introduce EchoJEPA, a foundation model demonstrating that latent prediction outperforms pixel reconstruction for echocardiography. Trained on 18 million videos, EchoJEPA reduces LVEF error by 27% over compute-matched baselines, achieves 78.6% view accuracy with 1% of labels (vs. 42.1% at 100%), degrades only 2.3% under acoustic perturbations (vs. 16.8%), and transfers zero-shot to pediatric patients better than fine-tuned models. We release EchoJEPA-L and our evaluation framework at https://github.com/bowang-lab/EchoJEPA.